# Finite-Time Analysis of Simultaneous Double Q-learning

### v1

@@ B.2 Convergence of $Q_{k}^{\text{err}_{UL}}$
The system (B.3) is a stochastic linear system with system matrix $I+\alpha\gamma DP\Pi_{Q_{k}^{*}}-\alpha D$ and noise $w_{k}^{A}-w_{k}^{B}$. To establish the convergence bound of this system, the same analysis approach as in Section A can be applied.
In particular, let us define $x_{k}\coloneqq Q_{k}^{\text{err}_{UL}}$ and $A\coloneqq I+\alpha\gamma DP\Pi_{Q_{k}^{*}}-\alpha D$. Then, the system (B.3) can be presented as the following stochastic linear system:
@@
$w_{k}^{A}-w_{k}^{B}=(e_{a_{k}}\otimes e_{s_{k}})(\delta_{k}^{A}-\delta_{k}^{B})-\gamma DP(\Pi_{Q^{B}_{k}}Q^{A}_{k}-\Pi_{Q^{A}_{k}}Q^{B}_{k})+D(Q_{k}^{A}-Q_{k}^{B}),\$
@@
To prove the convergence of $Q_{k}^{\text{err}_{UL}}$, we prove the boundedness of the noise term in (20).
The boundedness of $w_{k}^{A}-w_{k}^{B}$ in (20) is formally proved in the following lemma.
@@
$\mathbb{E}\left[(w_{k}^{A}-w_{k}^{B})^{T}(w_{k}^{A}-w_{k}^{B})\right]\leq\frac{16}{(1-\gamma)^{2}}\coloneqq W_{\text{max}}$
@@
$\begin{split}&\mathbb{E}\left[(w_{k}^{A}-w_{k}^{B})^{T}(w_{k}^{A}-w_{k}^{B})\right]\\ =&\mathbb{E}\left[\lVert((e_{a_{k}}\otimes e_{s_{k}})(\delta_{k}^{A}-\delta_{k}^{B}))-(\gamma DP(\Pi_{Q^{B}_{k}}Q^{A}_{k}-\Pi_{Q^{A}_{k}}Q^{B}_{k})-D(Q_{k}^{A}-Q_{k}^{B}))\rVert_{2}^{2}\right]\\ =&\mathbb{E}\left[(\delta_{k}^{A}-\delta_{k}^{B})^{2}\right]-\lVert\gamma DP(\Pi_{Q^{B}_{k}}Q^{A}_{k}-\Pi_{Q^{A}_{k}}Q^{B}_{k})-D(Q_{k}^{A}-Q_{k}^{B}))\rVert_{2}^{2}\\ \leq&\mathbb{E}\left[(\delta_{k}^{A}-\delta_{k}^{B})^{2}\right]\\ =&\mathbb{E}\left[(r_{k}^{A}+\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{B}}Q_{k}^{A}-(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}-(r_{k}^{B}+\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{A}}Q_{k}^{B}-(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}))^{2}\right]\\ \leq&\mathbb{E}\left[(|r_{k}^{A}|+|\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{B}}Q_{k}^{A}|+|(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}|+(|r_{k}^{B}|+|\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{A}}Q_{k}^{B}|+|(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}|))^{2}\right]\\ =&\frac{16}{(1-\gamma)^{2}}=W_{\text{max}}\end{split}$
Now, because (B.3) is a stochastic linear system, the analysis of a simple stochastic linear system from Section A can be applied directly. Then, we can get the upper bound of $Q_{k}^{\text{err}_{UL}}$ in the following lemma.
@@
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{2}]\leq\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\rho^{k}$
@@
$\begin{split}\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{2}^{2}]&=\mathbb{E}[\lVert(Q_{k}^{\text{err}_{UL}})^{T}(Q_{k}^{\text{err}_{UL}})\rVert]\\ &=\mathbb{E}[\lVert\text{tr}(Q_{k}^{\text{err}_{UL}})^{T}(Q_{k}^{err_{UL}})\rVert]\\ &=\mathbb{E}[\lVert\text{tr}((Q_{k}^{\text{err}_{UL}})(Q_{k}^{err_{UL}})^{T}\rVert]\\ &=\mathbb{E}[\lVert\text{tr}(X_{k})\rVert]\end{split}$
@@
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{2}^{2}]\leq\frac{16\alpha\lvert S\times A\rvert^{2}}{d_{\text{min}}(1-\gamma)^{3}}+\lvert S\times A\rvert^{2}\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}^{2}\rho^{2k}$
@@
$=(I-\alpha D)(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}})$
$\quad+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{UL}}+\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{UL}}-\Pi_{Q_{k}^{*}}Q_{k}^{\text{err}_{UL}})$
$=(I-\alpha D+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}})(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}})$
$\quad+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}-\Pi_{Q_{k}^{*}})Q_{k}^{\text{err}_{UL}}$ (21)
@@
$\leq\rho\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$
$\quad+\alpha\gamma d_{\text{max}}\biggl{(}\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\rho^{k}\biggl{)},$ (22)
@@
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$ $\leq\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$ (23)
Using this result, we can obtain the bound of $\mathbb{E}\bigl{[}\lVert Q_{k}^{\text{err}_{U}}\rVert_{\infty}\bigr{]}$.
Thus, $Q_{k}^{\text{err}_{U}}$ satisfies
@@
$\leq\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\rho^{k}$
$\leq\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k},$ (24)
where the second equality is due to (23) and Lemma 3 and the last inequality is due to the following fact $\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\leq\lvert S\times A\rvert^{1/2}\lVert Q_{0}^{\text{err}_{UL}}\rVert_{\infty}\leq\lvert S\times A\rvert^{1/2}\frac{2}{1-\gamma}$. Because the upper comparison system bounds all trajectory that of original system, we use this bound as the upper bound of the original system.
B.3 CONVERGENCE OF $Q_{k}^{\text{err}_{L}}$
@@
$Q_{k+1}^{\text{err}_{L}}=(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)Q_{k}^{\text{err}_{L}}+\alpha w_{k}^{A}-\alpha w_{k}^{B},\quad Q_{0}^{\text{err}_{L}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$
@@
Proposition B.3.
Suppose $Q_{0}^{\text{err}_{L}}\leq Q_{0}^{\text{err}}$, where $\leq$” is used as the element-wise inequality. Then, we have
$Q_{k}^{\text{err}_{L}}\leq Q_{k}^{\text{err}}$
@@
$=(I-\alpha D)(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}})$
$\quad+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{{B}}}Q_{k}^{\text{err}_{L}}+\Pi_{Q_{k}^{{B}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{B}}Q_{k}^{\text{err}_{U}})$
@@
$\leq\rho\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}\rVert_{\infty}]$
$\quad+\alpha\gamma d_{\text{max}}\biggl{(}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}\biggl{)},$ (26)
@@
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}\rVert_{\infty}]$ $\leq\rho^{k}\mathbb{E}[\lVert Q_{0}^{\text{err}_{U}}-Q_{0}^{\text{err}_{L}}\rVert_{\infty}]+\alpha\gamma d_{\text{max}}\biggl{(}\frac{\rho^{k}}{1-\rho}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+\frac{\rho^{k}}{1-\rho}\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\rho^{k-2}\frac{(k-1)(k-2)}{2}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+k\rho^{k-1}\frac{2\lvert S\times A\rvert^{3/2}}{1-\gamma}\biggl{)}$
$=\rho^{k}\frac{4\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\rho^{k}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+\rho^{k-2}(k-1)(k-2)\frac{\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+k\rho^{k-1}\frac{2\lvert S\times A\rvert^{3/2}\alpha\gamma d_{\text{max}}}{1-\gamma},$ (27)

### v2

@@ B.2 Convergence of $Q_{k}^{\mathrm{err}_{UL}}$
The system (B.3) is a stochastic linear system with system matrix $I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D$ and noise $w_{k}^{A}-w_{k}^{B}$. To establish the convergence bound of this system, the same analysis approach as in A can be applied. In particular, let us define $x_{k}\coloneqq Q_{k}^{\text{err}_{UL}}$ and $A\coloneqq I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D$. Then, the system (B.3) can be presented as the following stochastic linear system:
@@
$w_{k}^{A}-w_{k}^{B}=(e_{a_{k}}\otimes e_{s_{k}})(\delta_{k}^{A}-\delta_{k}^{B})-\gamma DP(\Pi_{Q^{B}_{k}}Q^{A}_{k}-\Pi_{Q^{A}_{k}}Q^{B}_{k})+D(Q_{k}^{A}-Q_{k}^{B}),\\$
@@
Here, $r_{k+1}^{A}$ and $r_{k+1}^{B}$ denote the instantaneous rewards observed at iteration $k$ for the updates of $Q_{k}^{A}$ and $Q_{k}^{B}$, respectively. To prove the convergence of $Q_{k}^{\text{err}_{UL}}$, we prove the boundedness of the noise term in (21). The boundedness of $w_{k}^{A}-w_{k}^{B}$ in (21) is formally proved in the following lemma.
@@
$\mathbb{E}\left[(w_{k}^{A}-w_{k}^{B})^{T}(w_{k}^{A}-w_{k}^{B})\right]\leq\frac{16}{(1-\gamma)^{2}}\coloneqq W_{\mathrm{max}},$
@@
$\mathbb{E}\bigl[(w_{k}^{A}-w_{k}^{B})^{T}(w_{k}^{A}-w_{k}^{B})\bigr]$
$=\mathbb{E}\Bigl[\bigl\lVert(e_{a_{k}}\otimes e_{s_{k}})(\delta_{k}^{A}-\delta_{k}^{B})-\bigl(\gamma DP(\Pi_{Q_{k}^{B}}Q_{k}^{A}-\Pi_{Q_{k}^{A}}Q_{k}^{B})-D(Q_{k}^{A}-Q_{k}^{B})\bigr)\bigr\rVert_{2}^{2}\Bigr]$
$=\mathbb{E}\bigl[(\delta_{k}^{A}-\delta_{k}^{B})^{2}\bigr]-\Bigl\lVert\gamma DP(\Pi_{Q_{k}^{B}}Q_{k}^{A}-\Pi_{Q_{k}^{A}}Q_{k}^{B})-D(Q_{k}^{A}-Q_{k}^{B})\Bigr\rVert_{2}^{2}$
$\leq\mathbb{E}\bigl[(\delta_{k}^{A}-\delta_{k}^{B})^{2}\bigr]$
$=\mathbb{E}\Bigl[\bigl(r_{k+1}^{A}+\gamma\,e_{s_{k}^{\prime}}^{T}\Pi_{Q_{k}^{B}}Q_{k}^{A}-(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}-\,\bigl(r_{k+1}^{B}+\gamma\,e_{s_{k}^{\prime}}^{T}\Pi_{Q_{k}^{A}}Q_{k}^{B}-(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}\bigr)\bigr)^{2}\Bigr]$
$\leq\mathbb{E}\Bigl[\bigl(\lvert r_{k+1}^{A}\rvert+\lvert\gamma\,e_{s_{k}^{\prime}}^{T}\Pi_{Q_{k}^{B}}Q_{k}^{A}\rvert+\lvert(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}\rvert+\,\lvert r_{k+1}^{B}\rvert+\lvert\gamma\,e_{s_{k}^{\prime}}^{T}\Pi_{Q_{k}^{A}}Q_{k}^{B}\rvert+\lvert(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}\rvert\bigr)^{2}\Bigr]$
$=\frac{16}{(1-\gamma)^{2}}=W_{\max}$
This bound on the noise term plays a key role in establishing the finite-time convergence of the stochastic linear system (21). Now, because (B.3) is a stochastic linear system, the analysis of a simple stochastic linear system from A can be applied directly. Then, we can get the upper bound of $Q_{k}^{\text{err}_{UL}}$ in the following lemma.
@@
$\mathbb{E}[\lVert Q_{k}^{\mathrm{err}_{UL}}\rVert_{2}]\leq\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\mathrm{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert\mathcal{S}\times\mathcal{A}\rvert\lVert Q_{0}^{\mathrm{err}_{UL}}\rVert_{2}\rho^{k}.$
@@
$\begin{split}\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{2}^{2}]&=\mathbb{E}[(Q_{k}^{\text{err}_{UL}})^{T}(Q_{k}^{\text{err}_{UL}})]\\ &=\mathbb{E}[\text{tr}(Q_{k}^{\text{err}_{UL}})^{T}(Q_{k}^{err_{UL}})]\\ &=\mathbb{E}[\text{tr}((Q_{k}^{\text{err}_{UL}})(Q_{k}^{err_{UL}})^{T}]\\ &=\mathbb{E}[\text{tr}(X_{k})]\end{split}$
@@
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{2}^{2}]\leq\frac{16\alpha\lvert\mathcal{S}\times\mathcal{A}\rvert^{2}}{d_{\text{min}}(1-\gamma)^{3}}+\lvert\mathcal{S}\times\mathcal{A}\rvert^{2}\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}^{2}\rho^{2k}$
Taking the square root on both side of the last inequality, using the subadditivity of the square root function, the Jensen inequality, and the concavity of the square root function, we have the desired conclusion.
@@
$=(I-\alpha D)(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}})+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{UL}}$
$\quad+\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{UL}}-\Pi_{Q_{k}^{*}}Q_{k}^{\text{err}_{UL}})$
$=(I-\alpha D+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}})(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}})+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}-\Pi_{Q_{k}^{*}})Q_{k}^{\text{err}_{UL}}$ (22)
@@
$\leq\rho\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]+\alpha\gamma d_{\text{max}}\biggl(\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert\mathcal{S}\times\mathcal{A}\rvert\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\rho^{k}\biggl),$ (23)
@@
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$ $\leq\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}.$ (24)
Using this result, we can obtain the bound of $\mathbb{E}\bigl[\lVert Q_{k}^{\text{err}_{U}}\rVert_{\infty}\bigr]$. Thus, $Q_{k}^{\text{err}_{U}}$ satisfies
@@
$\leq\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}+\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert\mathcal{S}\times\mathcal{A}\rvert\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\rho^{k}$
$\leq\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}+\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k},$ (25)
where the second equality is due to (24) and Lemma 7 and the last inequality is due to the following fact $\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\leq\lvert\mathcal{S}\times\mathcal{A}\rvert^{1/2}\lVert Q_{0}^{\text{err}_{UL}}\rVert_{\infty}\leq\lvert\mathcal{S}\times\mathcal{A}\rvert^{1/2}\frac{2}{1-\gamma}$. Because the upper comparison system bounds all trajectory that of original system, we use this bound as the upper bound of the original system.
B.3 CONVERGENCE OF $Q_{k}^{\mathrm{err}_{L}}$
@@
$Q_{k+1}^{\text{err}_{L}}=(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)Q_{k}^{\text{err}_{L}}+\alpha w_{k}^{A}-\alpha w_{k}^{B},\quad Q_{0}^{\text{err}_{L}}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$,
@@
Proposition A.3.
Suppose $Q_{0}^{\mathrm{err}_{L}}\leq Q_{0}^{\mathrm{err}}$, where $\leq$” is used as the element-wise inequality. Then, we have
$Q_{k}^{\mathrm{err}_{L}}\leq Q_{k}^{\mathrm{err}},$
@@
$=(I-\alpha D)(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}})+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{{B}}}Q_{k}^{\text{err}_{L}}+\Pi_{Q_{k}^{{B}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{B}}Q_{k}^{\text{err}_{U}})$
@@
$\leq\rho\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}\rVert_{\infty}]+\alpha\gamma d_{\text{max}}\biggl(\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}\biggl),$ (27)
@@
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}\rVert_{\infty}]$ $\leq\rho^{k}\mathbb{E}[\lVert Q_{0}^{\text{err}_{U}}-Q_{0}^{\text{err}_{L}}\rVert_{\infty}]+\alpha\gamma d_{\text{max}}\biggl(\frac{\rho^{k}}{1-\rho}\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{\rho^{k}}{1-\rho}\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}$
$\quad+\rho^{k-2}\frac{(k-1)(k-2)}{2}2\alpha\gamma d_{\text{max}}\frac{\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}+k\rho^{k-1}\frac{2\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}\biggl)$
$=\rho^{k}\frac{4\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\rho^{k}\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\rho^{k-2}(k-1)(k-2)\frac{\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}$
$\quad+k\rho^{k-1}\frac{2\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}\alpha\gamma d_{\text{max}}}{1-\gamma},$ (28)